C.M. is an 89-year-old female who presents with her daughter to the provider’s office. The daughter is concerned because her mother has not been eating well.

SUBJECTIVE DATA

 PMH: HTN, Hypothyroid  24-hour diet recall reveals approximately 1,100-calorie intake  Lives alone, daughter nearby  Does not drive  Daughter shops  Patient cooks light meals

OBJECTIVE DATA  Vital signs: T 37 P 72 R 12 BP 104/64  Weight: 105 lb  Weight last visit, 3 months ago: 115 lb  HT: 5 feet, 1 inch  Medications: Metoprolol 12.5 mg per day, Synthroid 0.75 mg per day  Oral mucosa dry, gums bleeding, dentures loose  Poor skin turgor Respond to each question for the following scenario. Use rationale to support your responses.

1. Develop a problems list from the objective and subjective data provided. Include at least three (3) problems with rationale.

2. What is the definition of body mass index, and how is it calculated? With the information provided, what is the patient’s body mass index?

3. How many calories a day does this patient need to maintain her current weight? How is this determined?

4. What laboratory studies would help determine the patient’s nutritional status and general overall health? List and describe at least three laboratory studies with rationale.
Answer #1

1. Objective problem list :-

- weight loss because of not eating well

- poor skin turgor and oral mucosa dryness because less intake of fluids and electrolytes

- loose dentures due to increased age

Subjective problems list :-

- she has Hypertension and hypothyroidism

2. Body mass index describes relative weight for height .Body Mass Index (BMI) is a person's weight in kilograms divided by the square of height in meters. A high BMI can be an indicator of high body fatness. BMI can be used to screen for weight categories that may lead to health problems .

3. Calories needed by this patient is :

Age is 89 years

Sedentary lifestyle

Weight is 105lbs = 105×0.45kg = 47.2 kg

Height is 5feet 1inch = 154.9cm

BMR = 10W + 6.25H -5A-161

= 10×47.2 + 6.25× 154.9 - 5×89 - 161

= 472+ 968.1 - 445 - 161

= 1440.1 - 445 - 161

= 834.1

For sedentary lifestyle , calories = BMR × 1.2 = 1000.92 calories

4. Laboratory studies include :-

- LFT to check protein level

- CBC to check Hemoglobin levels

- RFT to check any toxicities